Circular No. 8229

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 U2 (LINEAR) F. Shelly, Lincoln Laboratory, Massachusetts Institute of Technology,

reports the discovery of a comet with a diffuse coma and a very wide, fanshaped tail in p.a. 85^{o} on LINEAR images (discovery observation below). Following posting on the NEO Confirmation Page, other CCD observers have also commented on the cometary appearance on Oct. 21.1–21.2 UT, including J. Young at Table Mountain (0.6-m reflector; 5′′ coma without central condensation and with a fan-shaped tail ∼ 25′′ long spanning p.a. 70^{o}–95^{o}) and R. Fredrick and T. Medlock at Louisburg, KS (0.75-m reflector; 30′′ tail in p.a. 80^{o}).

2003 UT  α_{2000}  δ_{2000}  Mag.
Oct. 19.08914  20^{h}19^{m}13.^{s}43  −3^{o}35′46′′.3  18.3

The available astrometry (including prediscovery LINEAR observations on Sept. 19 and Oct. 2), the following orbital elements, and an ephemeris appear on MPEC 2003-U38.

T = 2003 Dec. 4.9117 TT  ω = 177.^{o}3423
e = 0.621075  Ω = 186.4689 }2000.0
q = 1.710041 AU  i = 24.4852
a = 4.512875 AU n^{o} = 0.1028073 P = 9.6 years

SUPERNOVA 2003ix IN UGC 3746 The Lick Observatory Supernova Search program (cf. IAUC 6627),

which is led by A. V. Filippenko and W. Li (and currently includes B. Beutler, J. Burket, M. Ganeshalingam, J. Graham, H. Pugh, M. Moore, D. Singer, B. Swift, and D. Weisz), reports (via Li) the discovery by Ganeshalingam of an apparent supernova (mag ∼ 18.8) on unfiltered KAIT images taken on Oct. 19.5 and 20.5 UT. The new object is located at α = 7^{h}14^{m}25^{s}.72, δ = +55^{o}08′53′′.4 (equinox 2000.0), which is 7′′.0 west and 1′′.1 south of the nucleus of UGC 3746. A KAIT image taken on Apr. 7.1 showed nothing at this position (limiting mag ∼ 19.5).

DE CIRCINI Additional magnitudes by W. Liller (cf. IAUC 8221): Oct. 13.000 UT,

V = 10.76; 13.004, B = 11.12; 16.999, V = 12.45; 17.995, V = 12.60.

2003 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green